OpenStudy (anonymous):

Given A (-6,4) and M(-1, -1) find B.

OpenStudy (anonymous):

@telijahmed can u please help me?

OpenStudy (anonymous):

is M the midpoint

OpenStudy (anonymous):

if m is the midpoint then it means \[\frac{ -6+x }{ 2 }=-1\]

OpenStudy (anonymous):

where x is the value of the x value of B

OpenStudy (anonymous):

\[\frac{ 4+y }{ 2}=-1\]

OpenStudy (anonymous):

if you gey x and y then those are the values of B

OpenStudy (anonymous):

IS THAT OK @xolasia